OpenStudy (anonymous):

please help me! 1. Is x = –3 a solution to the equation –3x2 = –27? 2. Is (–4, 10) a solution to the equation 10 – 5x = y? 3. Describe the pattern in the table using an equation. x/y chart 4. Solve. x + 5 = –2 5. Solve. 3b = 42 6. Solve. 7(s – 5) = 42 7. Solve. 10 = c over 3 – 4 + c over 6 8. Solve. 2/3x + 4 = 3/5x – 2 9. Solve. 3(h – 4) = –1/2(24 – 6h) 10. Solve for x. ax + bx = –c 11. You work for 4 hours on Saturday and 8 hours on Sunday. You also receive a $50 bonus. You earn

OpenStudy (anonymous):

Hm. You need to pick one. 1. I don't know what you're trying to say. 2. Plug in your x and y and see if both sides are the same. 3. You copy/pasted. There's no chart. 4. Subtract 5 from each side to get x by itself. 5. Divide both sides by 3 to get b by itself. 6. Divide both sides by 7, then add 3 to both sides to get s by itself. 7. Use the equation button down there to show us what it really looks like. Without parenthesis, it's really hard to read and understand what you mean. 8. Is that \[\frac{ 2 }{3x }+4=\frac{ 3 }{ 5x}+2\] or\[\frac{ 2 }{3 }x+4=\frac{ 3 }{ 5 }x+2\] 9. Multiply both sides by -2 to get rid of the fraction on the right. Pull out a common factor from what remains on the right. Divide both sides by that number. Distribute the constant you now have on the left through the parenthesis (hint: it will just change the signs of what's inside). You should now have a true statement. 10. Pull out the common factor from what's on the left. Divide both sides by what remains in the parenthesis to get x by itself. 11. ...

OpenStudy (anonymous):

1. yes it is a solution because when you plug it in -3(-3)^2=-27 -3(9)=-27 -27=-27 2. you want to plug in x and see if the y-value of the point is true so 10 – 5(-4) = y y=10+20 y=30 so no the point (-4,10) is not a solution 3. I need to see the chart 4. x+5=–2 subtract 5 from both sides x=-7 5.3b=42 divide both sides by 3 b=14 6.7(s – 5)=42 distribute 7 over the parantheses 7x-35=42 add 35 to both sides 7x=77 x=11

OpenStudy (anonymous):

7. can you use parantheses to make it more understandable 8.2/3x + 4 = 3/5x – 2 get the x values on one side but first get the denominator the same 10/15x+4=6/15x-2 subtract 6/15x from both sides 4/15x+4=-2 substract -4 from both sides 4/15x=-6 multiply both sides by 15/4 x=-45/2 or x=-22.5

OpenStudy (anonymous):

9. 3(h – 4)=–1/2(24–6h) distribute over the parantheses 3h-12=-12+3h 3h-12=3h-12 the equations are the same therefore h=all real numbers

OpenStudy (anonymous):

10. ax + bx = –c since there are no numbers all we can get is an expression (a+b)x=-c due to combining like terms we would but (a+b)x an examble is 2x+3x=5x which is (2+3)x if you break it down then divide both sides by a+b x=-c/(a+b)

OpenStudy (anonymous):

11. you need to finish it